# Wormhole Formation with Gravitational Particle Creation Mechanism

### v2

@@ Abstract
In this paper, we explore the possibility of the formation of a traversable wormhole geometry in General Relativity sustained by particle creation mechanism. Gravitational particle creation mechanism is a thermodynamic approach to generate repulsive back-reaction gravitational effects through the thermodynamic creation of particles. This repulsive back-reaction pressure can be thought of as a source of sustaining a traversable wormhole. In the first part of this paper, we systematically model a wormhole geometry through the process of particle creation by assuming an inverse square variation law of particle creation pressure within the wormhole geometry, and the shape function for the wormhole is obtained, assuming the redshift function as constant and finite. The obtained shape function is found to adhere to the feasibility conditions (flaring out condition and asymptotic flatness condition) of a viable wormhole geometry. In the second part of the paper, again by assuming a constant and finite redshift function, we studied the radial evolutionary behaviour of the particle creation pressures for three commonly utilized toy shape-functions. The examination of the energy conditions namely the Null Energy Condition (NEC) and the Dominant Energy Condition (DEC) for all these models reveal that the NEC is violated for the shape-function derived from the first principles of particle creation mechanism, as well as for the toy shape functions. The DEC is however found to be satisfied for all the aforementioned cases. In essence, NEC violation occurs through the particle creation phenomena inside the wormhole structure in spite of the existence of non-exotic matter.
@@ I Introduction
A class of physically reasonable traversable wormhole structures are first described by Morris and Thorne in GR to explore the possibility of interstellar space travel without violating the laws of special relativity (SR)[1]. Earlier studies demonstrated the possibility of traversable wormholes in GR [2, 3]. A more technical definition of a traversable wormhole is a topological rupture in the fabric of spacetime that allows for traversable interstellar connections as well as other temporal links between distant spacetime points. A traversable wormhole with exotic matter leads one to expect that the energy conditions are violated in the wormhole. For instance, it was argued that Casimir energy can be a possible source to sustain a traversable wormhole geometry [4, 5]. Several works validating the requirement of violation of the energy conditions for the existence of wormholes can be found in [6, 7, 8, 9, 10, 11, 12] .
@@
Observational data suggests that a significant portion of the total matter content is probably occupied by dark energy, which is responsible for the accelerated expansion of the Universe. As a basis of this observational result, the background dynamics are provided by the GR (as well as modified theories of gravity). Detailed information can be found in references [13, 14]. Generically, a cosmological fluid capable of generating negative pressure can be accounted as a possible explanation for the so-called dark energy. In the literature, there are several models of dark energy, which theoretically explain the accelerated expansion of the Universe with relevance to observed cosmological data [15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. In this context, a possible explanation for the observed accelerated expansion of the Universe, gravitational particle creation, is a theory based on non-equilibrium thermodynamics proposed by Prigogine [25], provides a mechanism capable of generating negative pressure in the Universe, through the production of particles involving thermodynamic processes at the expense of the gravitational field. Its primary advantage is that the negative pressure as a result of the creation of particles (also called back-reaction) is naturally accommodated by the second law of thermodynamics. Since the particle creation process is macroscopically described by a negative back-reaction pressure of quantum origin, these models have been studied as a possible means to explain the observed phenomena of the dark sector of cosmology [26, 27, 28]. Motivated by the aforementioned context, we suggest that the negative pressure produced as a result of the particle creation processes when incorporated into the energy-momentum tensor of the EFE, can be expected to contribute to the violation of the energy conditions required for the formation of wormholes. Therefore, we plan the paper in the following way: In section II, we develop the theoretical setting of the wormhole in the presence of gravitational particle creation and hence derive the shape function that describes the wormhole geometry. In section III, we assume different toy models of some viable wormhole geometry and obtain how particle creation pressure evolves inside the wormhole geometry. In section IV, we derive the embedding function and provide a 3D visualization of the wormhole shape function obtained through the requirements of particle creation. In section V, we finally discuss and conclude the results of our study.
II Wormhole Solution from Particle Creation Mechanism
The metric functions $\Phi(r)$ and $b(r)$ are general arbitrary functions of the radial coordinate $r$ within the context of a wormhole spacetime. The redshift function, denoted as $\Phi(r)$, characterizes the gravitational redshift effect, while the shape function, denoted as $b(r)$, provides information about the geometry of the wormhole. It is essential to note that the radial coordinate $r$ exhibits non-monotonic behavior, wherein it decreases from infinity to a minimum value $r_{0}$, representing the location of the throat of the wormhole. At the throat, $b(r_{0})=r_{0}$. Subsequently, the radial coordinate increases from $r_{0}$ to infinity.
[…]
In this paper, we address wormhole formation within the context of irreversible matter creation formalism and thermodynamics of open systems in cosmology, as described by Prigogine [25]. This formalism views the Universe as an open system, with particle formation described by an adaptation of the energy-momentum tensor that incorporates a matter creation term into conservation laws. Consider an open system of volume $V$, containing $N(t)$ particles with energy density $\rho$ and thermodynamic pressure $p$. The thermodynamic conservation equation in this system is
@@ II Wormhole Solution from Particle Creation Mechanism
In this equation, $dQ$ represents the heat absorbed by the system during time $dt$, $h=\rho+p$ is the enthalpy per unit volume, and $n=N/V$ is the particle density. In open systems, thermodynamic energy conservation includes the production and annihilation of matter, unlike isolated or closed systems where the number of particles is constant.
At this point, we use the formalism of irreversible matter creation from the thermodynamics of open systems to cosmology by considering a homogeneous and isotropic Universe as an open system with volume $V$, $N(t)$ particles, energy density $\rho$, and thermodynamic pressure $p$, as represented by the FLRW metric
$ds^{2}:=-dt^{2}+a^{2}(t)[dr^{2}+r^{2}d\Omega^{2}],\quad d\Omega^{2}=d\theta^{2}+sin^{2}\theta d\phi^{2}.$ (3)
In a homogeneous and isotropic Universe, the volume of the Universe may be represented using the scale factor $a$ as $V=a^{3}(t)$. Eq. (2) may be recast in terms of total time derivatives of physical quantities as follows [25]:
@@
In an open thermodynamical system, the non-conservation of the total rate of change of the number of particles, $N=na^{3}$, with comoving volume $a^{3}$ and $n$ being the number density of particle production, yields [29, 30]
@@
where $\Gamma$ is a particle creation rate that has negative and positive phases. Negative $\Gamma$ represents the particle destruction and positive $\Gamma$ describes the annihilation of particles.
The equation relating the creation rate $\Gamma$ to the creation pressure $\Pi$ is
@@
where $\Gamma$ is the source function, $\beta$ is an arbitrary constant in the range $[0,1]$ and $\omega$ is the barotopic equation of state (EoS) parameter. Here, as we are investigating the wormhole solution in the present era, we have set $H=H_{0}$, which is the present value of the Hubble parameter. Moreover, we assume that the matter inside the wormhole is perfect fluid. In this work, we consider specifically that the perfect fluid follows a dust-like equation of state, i.e $\omega=0$.
@@
The set of field equations necessary for our study then takes the form
@@
where $\rho(r)$ is the energy density, $p_{r}(r)$ and $p_{t}(r)$ are the radial and tangential pressures respectively which are mutually orthogonal. $\Pi(r)$ is the particle creation pressure introduced into the energy-momentum tensor $T^{\mu}_{\nu}=\text{diag}[-\rho(r),p_{r}(r)+\Pi(r),p_{t}(r)+\Pi(r),p_{t}(r)+\Pi(r)]$. Subsituting $\rho$ and (9) into Eq. (8) we obtain
@@
To obtain the explicit form of the shape function $b(r)$, the particle creation pressure should be specified in terms of the radial coordinate $r$. There may be different functional variations to how $\Pi(r)$ may evolve with $r$, but from a phenomenological standpoint, let us propose an inverse-square variation law of the particle creation pressure that varies with the radial coordinate $r$ as
@@
where $K$ is a proportionality constant, that accounts for the dimensions of $\Pi(r)$. This form of variation of particle creation pressure implies that the particle creation pressure is maximum at the throat ($r=r_{0}$) and decreases inverse-squarely with both radial directions. Nevertheless, one may still choose a general power-law variation ($\propto-1/r^{\alpha}$) as well. We considered the specific case by setting $\alpha=2$ for mathematical simplicity along to keep the model limited to the least number of free parameters.
Now, substituting Eq. (12) in Eq. (11), we obtain the shape function for our model as
@@
The necessary conditions for a viable wormhole geometry are
@@
    Throat condition: At the throat $r_{0}$, $b\left(r_{0}\right)=r_{0}$ and $01-\frac{b(r)}{r}>$ 0 outside the throat.
@@
    $\frac{b(r)-rb^{\prime}(r)}{b^{2}(r)}>0$
    , where ’ represents derivative w.r.t. $r$.
@@
Fig. 1 shows that the flareout and asymptotic conditions are fulfilled, with three sets of the parameter values $\beta$ and $K$. We can see clearly that the shape function (13) describes a viable wormhole geometry.

### v3

@@ Abstract
We explore the possibility of formation of a traversable wormhole in General Relativity supported by particle creation mechanism. The repulsive back-reaction pressure generated through this mechanism can be thought of as a source of sustaining a traversable wormhole. In the first part of this paper, we model a wormhole geometry by assuming an inverse powerlaw variation of particle creation pressure within the wormhole geometry, and the shape function for the wormhole is obtained, assuming a finite redshift function. By stabilizing the wormhole structure, based on the causality of sound-speed we obtained the existence range of the parameter $\beta$ associated with particle creation mechanism. The general shape function obtained is found to adhere to the feasibility conditions of a viable wormhole geometry. Then we studied the 2D and 3D embedding derived for the obtained shape function. In the second part of the paper, we followed the reverse approach of the aforementioned treatment, where we derived the particle creation pressures by assuming three commonly utilized toy shape-functions. The energy conditions are then investigated for all these cases. In essence, particle creation phenomena inside the wormhole may indeed provide for the possibility of sustenance of stable wormhole structure.
@@ I Introduction
A class of physically reasonable traversable wormhole structures are first described by Morris and Thorne in GR to explore the possibility of interstellar space travel without violating the laws of special relativity (SR)[1]. Earlier studies demonstrated the possibility of traversable wormholes in GR [2, 3]. A more technical definition of a traversable wormhole is a topological rupture in the fabric of spacetime that allows for traversable interstellar connections as well as other temporal links between distant spacetime points. A traversable wormhole with exotic matter leads one to expect that the energy conditions are violated in the wormhole. For instance, it was argued that Casimir energy can be a possible source to sustain a traversable wormhole geometry [4, 5]. Several works validating the requirement of violation of the energy conditions for the existence of wormholes can be found in [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. There are several important works, dealing with different frameworks, such as modified gravity and other physical coupling scenarios inspired from interaction with dark matter. Some of such work can be found in Refs. [17, 18, 19, 20, 21, 22, 23, 24, 25, 26].
@@
Observational data suggests that a significant portion of the total matter content is probably occupied by dark energy, which is responsible for the accelerated expansion of the Universe. As a basis of this observational result, the background dynamics are provided by the GR (as well as modified theories of gravity). Detailed information can be found in references [27, 28]. Generically, a cosmological fluid capable of generating negative pressure can be accounted as a possible explanation for the so-called dark energy. In the literature, there are several models of dark energy, which theoretically explain the accelerated expansion of the Universe with relevance to observed cosmological data [29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39]. In the context of accelerated expansion of the Universe, [40] proposed a theory of gravitational particle creation, which could be a competing candidate for explaining the observed accelerated expansion of the Universe and is based on non-equilibrium thermodynamics. This theory provides a mechanism capable of generating negative pressure in the Universe, through the production of particles involving thermodynamic processes at the expense of the gravitational field. Its primary advantage is that the negative pressure as a result of the creation of particles (also called back-reaction) is naturally accommodated by the second law of thermodynamics. Since the particle creation process is macroscopically described by a negative back-reaction pressure of quantum origin, these models have been studied as a possible means to explain the observed phenomena of the dark sector of cosmology [41, 42, 43]. Motivated by the aforementioned context, we suggest that the negative pressure produced as a result of the particle creation processes when incorporated into the energy-momentum tensor of the EFE, can be expected to contribute to the violation of the energy conditions required for the formation of wormholes. Therefore, we plan the paper in the following way: In section II, we briefly review the process of particle creation mechanism. In section III, we derive the shape function of a wormhole in the presence of particle creation mechanism. In section IV, we obtained the 2D and 3D embedding visualization of the wormhole shape function obtained through the requirements of particle creation. In section V, we derived the particle creation pressures inside the wormhole using various toy shape functions. In section VI, we finally discuss and conclude the results of our study.
II Brief Review of Gravitational Particle Creation
Particle creation models have been thoroughly investigated as a promising framework for the purpose of describing the cosmological dark sector. This is achieved through matter creation process characterized by negative pressure of quantum origin, known as backreaction, which was mentioned in the previous section. Within this context, the accelerated expansion of the Universe in all eras, including the early time inflation as well as expansion of the Universe in the late-time era, emerges as a result of the mechanism of gravitational particle production. Some of the earlier development of this interesting possibility can be found in Refs. [44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56]. Cosmologically inspired wormholes are interesting and have been studied in various cosmological context [57, 58, 59, 60, 61, 62]. Quite recently, evolving wormholes in a Friedmann Universe have been studied by Bronnikov [63]. Motivated by this likelihood of wormholes in various cosmological context, and relying on the assumption that a standard local wormhole spacetime can coexist locally in a global FLRW Universe, we intend to investigate the possibility of wormhole formation in a Friedmann Universe, within the background of gravitational particle creation mechanism.
To formulate the necessary equations of particle creation, one starts with the conservation equation of an open thermodynamic system of volume $V$ containing $N(t)$ number of particles, having energy density $\rho$ and thermodynamic pressure $p$. The conservation equation is
@@ II Brief Review of Gravitational Particle Creation
with $dQ$ as the heat recieved by the sytem in time $dt$. Here $h=\rho+p$ is the enthalpy per unit volume and $n=N/V$ is the number density of particles. One may find that, in contrast to closed systems, where the particle number stays constant, the thermodynamic energy conservation in open systems incorporates a term (third term in Eq.(1)) that represents the processes of matter creation and annihilation that occur within the system unlike isolated or closed systems where the number of particles is constant. The second law of thermodynamics takes the following form
$dS=d_{e}S+d_{i}S\geq 0,$ (2)
$d_{e}S$ denotes entropy flow, whereas $d_{i}S$ represents entropy generation. The first term of the equation reflects the system’s homogeneity, while the latter relates to the entropy associated with particle creation. In order to obtain expressions for these aforementioned quantities, one may start by expressing the total differential of the entropy given by
$\mathfrak{T}dS=d(\rho V)+pdV-\mu d(nV)$ (3)
with $\mu$ as the chemical potential and $\mathfrak{T}$ being the thermodynamic temperature. Now, by utilizing Eq. (1) and the standard relation $\mu n=h-\mathfrak{T}s$, where $s=S/V$ being the entropy density, it is possible to rewrite Eq. (3) in the form
$\mathfrak{T}dS=dQ+\mathfrak{T}\frac{s}{n}d(nV).$ (4)
From Eq. (2), we should have
$\mathfrak{T}dS=\mathfrak{T}d_{e}S+\mathfrak{T}d_{i}S,$ (5)
Using Eq. (5), one can obtain direct relations for entropy flow and entropy creation as
$d_{s}S=\frac{dQ}{\mathfrak{T}},\quad d_{i}S=\frac{s}{n}d(nV).$ (6)
Since the entropy flow may be considered as a measure of the variation in a system’s homogeneity, a homogeneous system exhibits no variation in homogeneity. As an outcome, the entropy flow is zero in such a system. It suggests that in homogeneous systems, adiabatic processes are expected to occur, and hence matter formation is the primary source of entropy generation, i.e.
$dS=d_{i}S=\frac{s}{n}d(nV)\geq 0.$ (7)
This mechanism of irreversible creation of particles can be applied to cosmology. In a homogeneous and isotropic Universe, the volume of the Universe may be represented using the scale factor $a$ as $V=a^{3}(t)$. Eq. (LABEL:pceq1) may be recast in terms of total time derivatives of physical quantities as follows [40]:
@@
Eq. (9) can be recast into a convenient form
$\dot{\rho}+3H(\rho+p+\Pi)=0,$ (10)
where
$\Pi=-\frac{\rho+p}{3Hn}(\dot{n}+3Hn)$ (11)
is the particle creation pressure. One may see that, the continuity Eq. (10) is a modified form of the usual continuity equation in GR, $\dot{\rho}+3H(\rho+p)=0$, where in an extra term $\Pi$ has been introduced as a consequence of matter creation. In an open thermodynamical system, the non-conservation of the total rate of change of the number of particles, $N=na^{3}$, with comoving volume $a^{3}$ and $n$ being the number density of particle production, yields [44, 45]
@@
where $\Gamma$ is a particle creation rate (or source function) that can have negative or positive signs. Negative sign of $\Gamma$ represents the particle annihilation and positive sign of $\Gamma$ describes particle creation. Comparing Eqs. (11) and (12), the equation relating the creation rate $\Gamma$ to the creation pressure $\Pi$ can be found as
@@
where $\omega$ is the barotopic equation of state (EoS) parameter that relates the energy density with pressure, i.e. $p=\omega\rho$. Here, as we are investigating the wormhole solution in the present era, we have set $H=H_{0}$, which is the present value of the Hubble parameter. Moreover, we assume that the matter inside the wormhole is perfect fluid. In this work, we consider specifically that the perfect fluid follows a dust-like equation of state, i.e $\omega=0$.
III Wormhole Solution from Particle Creation Mechanism
The metric functions $\Phi(r)$ and $b(r)$ are general functions of the radial coordinate $r$ within the context of a wormhole spacetime. The redshift function, denoted as $\Phi(r)$, characterizes the gravitational redshift effect, while the shape function, denoted as $b(r)$, provides information about the geometry of the wormhole. It is essential to note that the radial coordinate $r$ exhibits non-monotonic behavior, wherein it decreases from infinity to a minimum value $r_{0}$, representing the location of the throat of the wormhole. At the throat, $b(r_{0})=r_{0}$. Subsequently, the radial coordinate increases from $r_{0}$ to infinity.
[…]
In this paper, we address wormhole formation within the context of irreversible matter creation formalism and thermodynamics of open systems in cosmology, as described by Prigogine [40]. This formalism views the Universe as an open system, with particle formation described by an adaptation of the energy-momentum tensor that incorporates a matter creation term into conservation laws. Consider an open system of volume $V$, containing $N(t)$ particles with energy density $\rho$ and thermodynamic pressure $p$.
@@ III Wormhole Solution from Particle Creation Mechanism
where $\beta$ is an arbitrary constant of the order of unity. According to the authors of [48], the values of $\beta<<1$ lead to negligible particle creation effects and may be neglected. The set of field equations can be obtained by using the metric (15) in the Einstein field equations $G_{\mu\nu}=8\pi T_{\mu\nu}$ as
@@
where $\rho(r)$ is the energy density, $p_{r}(r)$ and $p_{t}(r)$ are the radial and tangential pressures respectively which are mutually orthogonal. $\Pi(r)$ is the particle creation pressure introduced into the energy-momentum tensor $T_{\mu\nu}=\text{diag}[-\rho(r),p_{r}(r)+\Pi(r),p_{t}(r)+\Pi(r),p_{t}(r)+\Pi(r)]$. Subsituting $\rho$ and (16) into Eq. (14) we obtain
@@
To obtain the explicit form of the shape function $b(r)$, the particle creation pressure should be specified in terms of the radial coordinate $r$. We argue that to sustain the wormhole structure the particle creation pressure needs to be larger at the throat (so that it can withstand the collapse of the throat) and then asymptotically decrease away from the throat. At the same time, the particle creation pressure should not be excessively large at the throat, as that would blow up the entire wormhole structure. Therefore, to obtain a sustained wormhole geometry, one must carefully chose a functional form of particle creation pressure $\Pi(r)$. For instance, if we assume the particle creation pressure in the form of inverse squared law like
@@
where $K$ is a proportionality constant that accounts for the dimensions of $\Pi(r)$, one finds the shape function in a linear form
@@
where we have substituted Eq. (19) in Eq. (18). Note that the linear shape function is a result of the inverse squared law given by Eq. (19), which is the simplest case. To have a more generalized picture, we may chose the inverse power-law form, i.e.,
$\Pi=-\frac{K}{r^{\alpha}},$ (21)
where $\alpha>0$ is a constant. The parameter $\alpha$, is tunable so that a non-linear form of shape function can be obtained. It is also seen that for Thus solving Eq. (18) with the help of Eq. (21), we get the general non-linear shape function
$b(r)=r_{0}-\frac{8\pi K\left(r^{3-\alpha}-r_{0}^{3-\alpha}\right)}{\beta(\alpha-3)(\omega+1)};\quad\alpha\neq 3,\omega\neq-1.$ (22)
The shape function given by Eq. (22) determines the configuration of the wormhole which is sustained by the negative particle creation pressure. The constraint on the possible values of $\alpha$ are $(0,3)\cup(3,\infty)$. However, to keep things simple, we choose to confine ourselves to the range $(0,3)$. In this work, we are motivated to investigate the possibility of a sustained wormhole structure in the presence of particle creation mechanism, without an external introduction of exotic fluids to account for the negative pressure required to sustain the wormhole geometry.
From Eq. (17), given the shape function (22), one may obtain the energy density, radial and tangential pressures given by
$\rho(r)=\frac{Kr^{-\alpha}}{\beta\omega+\beta},$ (23)
$p_{r}(r)=\frac{8\pi Kr^{-\alpha}}{(\alpha-3)\beta(\omega+1)}+Kr^{-\alpha}-\frac{8\pi Kr_{0}^{3-\alpha}}{(\alpha-3)\beta r^{3}(\omega+1)}-\frac{r_{0}}{r^{3}},$ (24)
$p_{t}(r)=\frac{r_{0}^{-\alpha}r^{-\alpha-3}\left(8\pi Kr_{0}^{3}r^{\alpha}-2Kr^{3}r_{0}^{\alpha}(4\pi(\alpha-2)-(\alpha-3)\beta(\omega+1))\right)}{2\beta(\alpha-3)(\omega+1)}+\frac{r_{0}}{2r^{3}},$ (25)
III.1 Constraints on $\beta$: Stability analysis
It is necessary to choose the parameter $\beta$ associated with particle creation in such a way that the obtained wormhole structure exhibits stability. In order to confirm the stability of the wormhole, we adopt the method of squared adiabatic sound speed given by [64]
$v_{s}^{2}=\frac{d\langle p\rangle}{dr}\left(\frac{d\rho}{dr}\right)^{-1}.$ (26)
where $\langle p\rangle=\frac{1}{3}(p_{r}+2p_{t})$, is the averaged pressure taken across the three spatial dimensions. This physical quantity represents the speed at which tiny acoustic perturbations or sound waves propagate through a particular medium during an adiabatic process. This allows us to probe into the stability of the wormhole in response to these small perturbations. The causality condition dictates $10\leq v_{s}^{2}<$ 1, which means that the speed limit of light cannot be exceeded by the perturbations. Using Eqs. (23), (24) and (25) we obtain
$v_{s}^{2}=\beta-\frac{8\pi}{3}$ (27)
To respect the causality condition we must have the following constraint on $\beta$
$\frac{8\pi}{3}\leq\beta<1+\frac{8\pi}{3},\quad\text{or}\quad 8.373\leq\beta<9.373.$ (28)
Having obtained the constrained values of $\beta$ within the range (28), we can now choose the values for $\beta$, that would ensure the stability of the wormhole. Let us now examine the validity of our derived shape function (22), through the viability conditions. The necessary conditions for a viable wormhole geometry are
@@ III.1 Constraints on $\beta$: Stability analysis
    Throat condition: At the throat $r_{0}$, $b\left(r_{0}\right)=r_{0}$ and $01-\frac{b(r)}{r}>$ 0 or $b(r)-r<0$ outside the throat.
@@
    $\frac{b(r)-rb^{\prime}(r)}{b^{2}(r)}>0,$
    where (’) represents derivative w.r.t. $r$.
@@
Fig. 1 shows the throat, flareout and asymptotic conditions are fulfilled, with three sets of the parameter values $\beta$ and $K$. The quantities $b^{\prime}(r)$, $b(r)/r$, $b(r)-r$ and $\frac{b(r)-rb^{\prime}(r)}{b^{2}(r)}$ are plotted for different combination of $\beta$, $K$ and $\alpha$. We can see clearly from the plots that the shape function (22) describes a viable wormhole geometry by adhering to the aforementioned conditions.
Figure 1: The viability conditions related to the shape function (22) is shown for three values of $K$, $\alpha$ and $\beta$
To explore the physical nature of the matter content inside the wormhole geometry, we shall make use of the energy conditions. A generic aspect of wormhole systems is the violation of the null energy condition (NEC) (i.e. $\rho+p_{r}\geq 0$ or $\rho+p_{t}\geq 0$ in terms of radial and tangential pressures respectively) at the throat of the wormhole structure. On the other hand, for some specific models keeping valid the NEC while keeping the dominant energy condition (DEC) ($\rho-p_{r}\geq 0$ or $\rho-p_{t}\geq 0$) violated at the throat is also possible [65, 66, 67, 68, 69]. In this work, we shall construct traversable wormhole structures with the violation of the NEC as well as DEC at the throat of the wormhole, simultaneously keeping the energy condition $\rho\geq 0$. This turns out, as a result of the violation of the weak energy condition (WEC) ($\rho\geq 0,\rho+p_{r}\geq 0\text{ or }\rho+p_{t}\geq 0$).
To study the energy conditions for the wormhole geometry (22), let us calculate the quantities $\rho+p_{r}$, $\rho+p_{t}$, $\rho-p_{r}$ and $\rho-p_{t}$ given by
$\rho+p_{r}:=\frac{K(\beta\omega+\beta+1)r^{-\alpha}}{\beta(\omega+1)}+\frac{8\pi K\left(r^{-\alpha}-\frac{r_{0}^{3-\alpha}}{r^{3}}\right)}{\beta(\alpha-3)(\omega+1)}-\frac{r_{0}}{r^{3}},$ (29)
$\rho+p_{t}:=\frac{r^{-\alpha-3}r_{0}^{-\alpha}\left(8\pi Kr_{0}^{3}r^{\alpha}+r_{0}^{\alpha}\left(2Kr^{3}((\alpha-3)(\beta\omega+\beta+1)-4\pi(\alpha-2))+(\alpha-3)\beta r_{0}(\omega+1)r^{\alpha}\right)\right)}{2\beta(\alpha-3)(\omega+1)},$ (30)
and
$\rho-p_{r}:=-\frac{K(\beta\omega+\beta-1)r^{-\alpha}}{\beta(\omega+1)}-\frac{8\pi K\left(r^{-\alpha}-\frac{r_{0}^{3-\alpha}}{r^{3}}\right)}{\beta(\alpha-3)(\omega+1)}+\frac{r_{0}}{r^{3}}.$ (31)
$\rho-p_{t}:=\frac{r^{-\alpha-3}r_{0}^{-\alpha}\left(r_{0}^{\alpha}\left(2Kr^{3}(4\pi(\alpha-2)-(\alpha-3)(\beta\omega+\beta-1))-(\alpha-3)\beta r_{0}(\omega+1)r^{\alpha}\right)-8\pi Kr_{0}^{3}r^{\alpha}\right)}{2\beta(\alpha-3)(\omega+1)}.$ (32)
Figure 2: The physical parameters and the energy conditions for the shape function (22) is shown for three combinations of $K$, $\alpha$ and $\beta$
Fig. 2 displays the behaviour of the related physical parameters $\rho(r),p_{r}(r),p_{t}(r)\text{ and }\omega_{r}=p_{r}(r)/\rho$ and the energy conditions like NEC and DEC for different combinations of the model parameters. As mentioned before, for the analysis we have set the EoS of the matter content inside the wormhole to follow a dust like $\omega=0$. Moreover, we shall choose the throat radius $r_{0}=1$. This choice is arbitrarily positive, provided the energy viability conditions are satisfied [70, 71]. The energy density as expected always takes positive values and is maximum at the throat and decreases asymptotically away from the throat. The radial pressure always takes negative values, which may be attributed to counteracting the radial collapse of the wormhole structure. Also, the tangential pressure always remain positive. The NEC in terms of radial pressure shows clear violation at the throat but obeys at far away regions from the throat, whereas the DEC is obeyed at the throat as well as at far away regions. However, in terms of the tangential pressure, the NEC is obeyed at the throat as well as in asymptotic regions but DEC is violated at the throat and satisfied at the asymptotic regions. Also as $\rho\geq 0$, one can say WEC is violated for the radial pressure case and obeyed for the tangential pressure one. To examine the behaviour of the effective matter resulting from the particle creation mechanism, we calculated the radial EoS parameter $\omega_{r}=p_{r}/\rho$, that essentially evolves with negative values while away from the throat it increases. This indicates that the effective matter at the throat behaves like exotic matter. In other words, we find that without explicitly introducing non-exotic matter, the wormhole is sustained due to the presence of particle creation mechanism in the interior of the wormhole structure. The negative pressure that arises at the throat is entirely due to particle creation mechanism.
IV Embedding
$ds^{2}=\left(1-\frac{b(r)}{r}\right)^{-1}dr^{2}+r^{2}d\phi^{2}.$ (33)
The metric (33) is equivalent to
$ds^{2}=dz^{2}+dr^{2}+r^{2}d\phi^{2},$ (34)
which can now be compared with Eq. (33) to obtain
$\frac{dz}{dr}=\pm\left(\frac{r}{b(r)}-1\right)^{-1/2},$ (36)
Eq. (36) can be solved to obtain a functional form of $z(r)$. However in our model given Eq. (22), the analytical form of a general $z(r)$ is not obtainable because of the complicated form of the shape function. Therefore, we adopt a numerical method to visualize the embedding function, which is shown in Fig. 3.
The embedding function obtained for the shape function (22), obtained from physical consideration of particle creation mechanism is shown in the Fig. 3 for $\beta=8.55$ and $K=0.005$.